OpenStudy (eva12):

Find the volume formed by rotating the region enclosed by:x=7y and y^3=x with y>0 about the -axis. Volume = ?

jhonyy9 (jhonyy9):

what is your opinion ,ideas ?

jhonyy9 (jhonyy9):

x=7y so for you graph it need adding values to x and calculi ,so when x = 0 you get y=0 too for x=1 result y = 1/7 for x=7 ---- y= 1 in case of second equation y^3 =x when y>0 for y = 1 result x= 1 -- y= 2 --- x= 8 so than now we graph these points

jhonyy9 (jhonyy9):